OpenStudy (hiyathere_20):

Write the equation of the quadratic function with roots 0 and 2 and a vertex at (1, 5).

OpenStudy (mhchen):

Isn't this technically the same problem as that last one? Roots 0 and 2 means if x=0 or 2, then y=0. y=(x-0)*(x-2) Vertex (1,5) means if x=1, then y=5. 5=a(1-0)*(1-2) 5=a*-1 a=-5 Answer: y=-5x(x-2) Which is expanded to \[y=-5x^2+10x\]